# Distillation-Free One-Step Diffusion for Real-World Image Super-Resolution

### v2

@@ Abstract
Diffusion models have been achieving excellent performance for real-world image super-resolution (Real-ISR) with considerable computational costs. Current approaches are trying to derive one-step diffusion models from multi-step counterparts through knowledge distillation. However, these methods incur substantial training costs and may constrain the performance of the student model by the teacher’s limitations. To tackle these issues, we propose DFOSD, a Distillation-Free One-Step Diffusion model. Specifically, we propose a noise-aware discriminator (NAD) to participate in adversarial training, further enhancing the authenticity of the generated content. Additionally, we improve the perceptual loss with edge-aware DISTS (EA-DISTS) to enhance the model’s ability to generate fine details. Our experiments demonstrate that, compared with previous diffusion-based methods requiring dozens or even hundreds of steps, our DFOSD attains comparable or even superior results in both quantitative metrics and qualitative evaluations. Our DFOSD also abtains higher performance and efficiency compared with other one-step diffusion methods. We will release code and models at https://github.com/JianzeLi-114/DFOSD.
@@ 1 Introduction
Real-world image super-resolution (Real-ISR) is a challenging task that aims to reconstruct high-resolution (HR) images from their low-resolution (LR) counterparts in real-world settings (Wang et al., 2020). Most image super-resolution (SR) methods (Kim et al., 2016; Johnson et al., 2016; Ledig et al., 2017; Chen et al., 2022; 2023) use Bicubic downsampling of HR images to generate LR samples for training and testing models. These methods achieve good results in reconstructing simple degraded images. However, they struggle with the complex and unknown degradations widely existing in real-world scenarios. Moreover, these methods often amplify the noise in LR images during reconstruction. Previous research has predominantly employed generative adversarial networks (GANs) (Goodfellow et al., 2020) architectures for image SR tasks (Wang et al., 2021a; Zhang et al., 2021; Liang et al., 2021). However, these approaches often struggle to train models that accurately capture real-world data distributions, leading to suboptimal generated content. Diffusion models (DMs), known for their strong denoising capabilities and ability to model data distributions, have been widely adopted in the field of image generation in recent years. Recently, numerous super-resolution (SR) methods based on pre-trained diffusion models have exhibited outstanding performance by leveraging their powerful priors and generative capabilities.
@@
Specifically, recent real-world image super-resolution (Real-ISR) models have predominantly leveraged powerful pre-trained diffusion models, such as large-scale text-to-image (T2I) models like Stable Diffusion (Wu et al., 2024b; Yang et al., 2024; Lin et al., 2024). With training on billions of image-text pairs and a strong capacity to model complex data distributions, these pre-trained T2I models provide extensive priors and powerful generative abilities. Most diffusion model (DM)-based methods generate high-resolution (HR) images by employing ControlNet models (Zhang et al., 2023), conditioning on the low-resolution (LR) inputs. However, these methods typically require tens to hundreds of diffusion steps to produce high-quality HR images. The introduction of ControlNet not only increases the number of model parameters but also further exacerbates inference latency. Consequently, DM-based multi-step diffusion methods often incur delays of tens of seconds when processing a single image, which significantly limits their practical application in real-world scenarios for low-level image reconstruction tasks, such as Real-ISR.
@@
To accelerate the generation process of diffusion models, recent research has introduced numerous one-step diffusion methods, known as diffusion distillation, which distill multi-step pre-trained diffusion models into one-step counterparts. Most of these approaches employ a knowledge distillation strategy, using the multi-step diffusion model as a teacher to train a one-step diffusion student model. These methods significantly reduce inference latency, and the quality of the generated images can be comparable to that of multi-step diffusion models. Real-ISR methods based on one-step diffusion models have become an increasingly popular research direction, with representative methods such as SinSR (Wang et al., 2024b) and OSEDiff (Wu et al., 2024a). While these methods achieve promising visual results, the inclusion of the teacher network increases training overhead. The performance of the student network is often constrained by the teacher network.
Refer to caption Canon_040 # Steps / MACs (G) Inference Time (s) Distillation-Free Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR StableSR DiffBIR ResShift SinSR OSEDiff DFOSD (ours) - / - 200 / 75,812 50 / 24,528 15 / 4,903 1 / 2,095 1 / 2,269 1 / 2,120 - 11.50 7.79 0.71 0.16 0.35 0.11 - N/A N/A N/A $\times$ $\times$ $\checkmark$
Figure 1: Visual comparisons ($\times$4) of different DM-based Real-ISR methods, including their inference times and MACs (Multiply-Accumulate Operations), for an output size of 512$\times$512. The inference times are measured on an A100 GPU. StableSR (Wang et al., 2024a), DiffBIR (Lin et al., 2024), and ResShift (Yue et al., 2024) are multi-step DM-based methods, performing 200, 50, and 15 sampling steps respectively. Our DFOSD is distillation-free when compared with other one-step diffusion models, like SinSR (Wang et al., 2024b) and OSEDiff (Wu et al., 2024a).
To overcome the aforementioned challenges, we propose DFOSD, a novel approach that generates HR images from their corresponding LR inputs in a single sampling step. Unlike previous one-step diffusion SR models, we do not employ knowledge distillation to train our one-step diffusion generator. Our approach eliminates the need to leverage outputs or corresponding noise from multi-step diffusion models, allowing us to train solely on real-world datasets. This significantly reduces training overhead and overcomes the limitations imposed by teacher models. Furthermore, we do not utilize models like CLIP (Radford et al., 2021) to encode prompts as conditional inputs for the diffusion model. Instead, we train a learnable text embedding. This further reduces the model’s inference time without compromising performance. As shown in Fig. 1, DFOSD not only achieves the best visual results but also attains the fastest inference speed.
To better leverage the prior knowledge of pre-trained multi-step models and enhance the authenticity of the generated images, we propose a noise-aware discriminator (NAD) initialized with parameters from the pre-trained stable diffusion (SD) UNet, which is trained adversarially alongside the generator. Specifically, our NAD takes the forward diffusion results of the latent features at various time steps, ensuring that its performance remains robust across different noise levels. NAD capitalizes on the prior knowledge of the pre-trained diffusion model, enhancing the reconstruction quality (see Fig. 1). Additionally, we propose edge-aware DISTS (EA-DISTS) loss to improve the authenticity of fine details in the generated content. Our comprehensive experiments indicate that DFOSD achieves superior performance and less inference time among one-step diffusion model (DM)-based Real-ISR models. When compared with multi-step DM-based models, DFOSD obtains comparable or even better performance with over 7$\times$ speedup in inference time (see Fig. 1).
@@
  • We propose DFOSD, a Distillation-Free One-Step Diffusion SR model training paradigm. Our DFOSD significantly enhances the details and visual quality of generated images, achieving remarkable results in both evaluation metrics and visual assessments.
  • We propose a noise-aware discriminator (NAD), which capitalizes on the prior knowledge from the pre-trained SD UNet and engages in adversarial training with the generator. Our NAD effectively enhances the realism and details of the reconstructed images.
  • We improve the perceptual loss used in image SR model training by proposing the edge-aware DISTS (EA-DISTS) loss. Our EA-DISTS leverages image edges to enhance the model’s ability and improve the authenticity of reconstructed details.
2 Related Works
@@ 2.1 Real-World Image Super-Resolution
Real-world image super-resolution (Real-ISR) aims to recover high-resolution (HR) images from low-resolution (LR) observations in real-world scenarios. The complex and unknown degradation patterns in such scenarios make Real-ISR a challenging problem (Ignatov et al., 2017; Liu et al., 2022a; Ji et al., 2020; Wei et al., 2020). To address this problem, models continuously evolve. Early image super-resolution models (Kim et al., 2016; Zhang et al., 2018c; b; Chen et al., 2022; 2023) typically rely on simple synthetic degradations like Bicubic downsampling for generating LR-HR pairs, resulting in subpar performance on real-world datasets. Later, GAN-based methods such as BSRGAN (Zhang et al., 2021), Real-ESRGAN (Wang et al., 2021a), and SwinIR-GAN (Liang et al., 2021) introduce more complex degradation processes. These methods achieve promising perceptual quality but encounter issues such as training instability. Additionally, they have limitations in preserving fine natural details. Recently, Stable Diffusion (SD) (Rombach et al., 2022b) is considered for addressing Real-ISR tasks due to its strong ability to capture complex data distributions and provide robust generative priors. Approaches such as StableSR (Wang et al., 2024a), DiffBIR (Lin et al., 2024), and SeeSR (Wu et al., 2024b) leverage pre-trained diffusion priors and ControlNet models (Zhang et al., 2023) to enhance HR image generation. While these methods significantly improve perceptual quality, the multi-step nature of diffusion models introduces latency issues, making them less practical for real-time applications in low-level image reconstruction tasks.
@@ 2.2 Acceleration of Diffusion Models
Acceleration of diffusion models can reduce computational costs and inference time. Therefore, various strategies have been developed to enhance the efficiency of diffusion models in image generation tasks. Fast diffusion samplers (Song et al., 2021; Karras et al., 2022; Liu et al., 2022b; Lu et al., 2022a; b; Zhao et al., 2024) have significantly reduced the number of sampling steps from 1,000 to 15$\sim$100 without requiring model retraining. However, further reducing the steps below 10 often leads to a performance drop. Under these circumstances, distillation techniques have made considerable progress in speeding up inference (Berthelot et al., 2023; Liu et al., 2022c; Meng et al., 2023; Salimans & Ho, 2022; Song et al., 2023; Zheng et al., 2023; Yin et al., 2024b; Liu et al., 2023; Geng et al., 2024). For instance, Progressive Distillation (PD) methods (Meng et al., 2023; Salimans & Ho, 2022) have distilled pre-trained diffusion models to under 10 steps. Consistency models (Song et al., 2023) have further reduced the steps to 2$\sim$4 with promising results. Instaflow (Liu et al., 2023) further achieves one-step generation through reflow (Liu et al., 2022c) and distillation. Recent score distillation-based methods, such as Distribution Matching Distillation (DMD) (Yin et al., 2024c; a) and Variational Score Distillation (VSD) (Wang et al., 2024c; Nguyen & Tran, 2024), aim to achieve one-step text-to-image generation. They minimize the Kullback–Leibler (KL) divergence between the generated data distribution and the real data distribution. Although these approaches have made notable progress, they still face challenges, like high training costs and dependence on teacher models.
3 Distillation-Free One-Step Diffusion (DFOSD)
Figure 2: Training framework of DFOSD. The left side represents the generator $\mathcal{G}_{\theta}$, which includes the pre-trained VAE and UNet from Stable Diffusion. Only the UNet is fine-tuned using LoRA, while other parameters remain frozen. The right side depicts the noise-aware discriminator (NAD), which guides the training process without participating in inference. The NAD extracts the UNet Mid-block outputs and processes them through an MLP to generate realism votes for different image regions. Both the downsampling and middle blocks of the UNet in the discriminator are fine-tuned with LoRA, whereas the MLP is randomly initialized.
In this section, we detail our Distillation-Free One-Step Diffusion (DFOSD) image super-resolution (SR) model. First, in Section 3.1, we review the fundamentals of diffusion models and introduce the principles underlying the DFOSD generator. Subsequently, we propose two key techniques for training our one-step diffusion SR model. In Section 3.2, we propose the noise-aware discriminator (NAD), which assesses image realism using the results of random forward diffusion applied to their latent representations. Then, in Section 3.3, we propose an improved perceptual loss function, edge-aware DISTS (EA-DISTS), designed to enhance the quality of image texture details. Finally, in Section 3.4, we outline the complete training process of the model.
3.1 Preliminaries: Diffusion
@@
As illustrated in Fig. 2, we first employ the encoder $E_{\theta}$ to map the low-resolution (LR) image $x_{L}$ into the latent space, yielding $z_{L}$: $z_{L}=E_{\theta}(x_{L})$. Next, we perform a single denoising step to obtain the predicted noise $\hat{\epsilon}$ and compute the high-resolution (HR) latent representation $\hat{z}_{H}$:
@@
where $\epsilon_{\theta}$ denotes the denoising network parameterized by $\theta$, and $T_{L}$ is the diffusion time step. Unlike one-step text-to-image (T2I) diffusion models (Song et al., 2023; Yin et al., 2024c), the input to the UNet of the Real-ISR diffusion models is not pure Gaussian noise. We set $T_{L}$ to an intermediate time step within the range [0, $T$], where $T$ is the total number of diffusion time steps. In Stable Diffusion (SD), $T=1,000$. Finally, we decode $\hat{z}_{H}$ using the decoder $D_{\theta}$ to reconstruct the HR image $\hat{x}_{H}$: $\hat{x}_{H}=D_{\theta}(\hat{z}_{H})$. The entire computation process of the generator can be expressed as $\hat{x}_{H}=\mathcal{G}_{\theta}(x_{L})$.
3.2 Noise-Aware Discriminator (NAD)
In an ideal scenario, we seek to achieve image restoration results that are almost indistinguishable from real images. Yet, training the generator directly without distillation often falls short of this goal. To improve the realism of generated images, we incorporate a discriminator. Training a discriminator from scratch, however, may result in unstable training dynamics, and converting the generator’s latent outputs to pixel space for evaluation introduces considerable computational overhead. Stable Diffusion (SD), a robust pre-trained generative model with strong priors and a UNet-based architecture, provides a promising solution to these challenges. This inspires us to initialize the discriminator with pre-trained UNet parameters, perform operations directly in the latent space, and leverage the UNet bottleneck layer’s robust information filtering and semantic condensation capabilities to construct the discriminator.
Figure 3 illustrates the visualization results of the latent representations of both generated and real images during the early stages of training. After undergoing forward diffusion at various random time steps, the UNet middle block outputs are visualized using dimensionality reduction techniques such as PCA, supervised UMAP, and LLE. The feature distributions of the generated images and real images exhibit distinct differences, thereby highlighting the Stable Diffusion (SD) UNet’s robust information filtering and semantic condensation capabilities.
Based on these observations, we propose a noise-aware discriminator (NAD). To better leverage the diffusion model’s ability to perceive noise at various levels and maintain the gap between generated and real images under different noise intensities, we feed the latent representations with randomly injected noise levels as inputs to the NAD. In Fig. 2, the NAD $\mathcal{D}_{\theta}$ consists of the UNet downsampling blocks (i.e., UNet Down-block in Fig. 2) and middle block (i.e., UNet Mid-block in Fig. 2), along with a MLP mapping the features into realism scores for different regions. $\mathcal{D}_{\theta}$ is initialized with the corresponding parameters of the SD UNet at the beginning of training. During training, we feed into the discriminator the forward diffusion results of both the latents predicted by the generator (i.e., $\hat{z}_{H}$ as mentioned in Eq. 1) and the corresponding ground truth latent vectors $z_{H}=E_{\theta}(x_{H})$.
Figure 3: Visualization of features dimensionality reduction for the first 100 channels from the middle block outputs of the Stable Diffusion (SD) UNet. Notably, there is a significant difference in the feature distributions at the UNet’s intermediate layers between real images and those generated by the one-step diffusion model during early training stages. This observation suggests that the intermediate layer features of the UNet are a robust basis for assessing image realism.
The adversarial losses for updating the generator and discriminator are defined as:

### v3

# Unleashing the Power of One-Step Diffusion based Image Super-Resolution via a Large-Scale Diffusion Discriminator

@@ Abstract
Diffusion models have demonstrated excellent performance for real-world image super-resolution (Real-ISR), albeit at high computational costs. Most existing methods are trying to derive one-step diffusion models from multi-step counterparts through knowledge distillation (KD) or variational score distillation (VSD). However, these methods are limited by the capabilities of the teacher model, especially if the teacher model itself is not sufficiently strong. To tackle these issues, we propose a new One-Step Diffusion model with a larger-scale Diffusion Discriminator for SR, called D3SR. Our discriminator is able to distill noisy features from any time step of diffusion models in the latent space. In this way, our diffusion discriminator breaks through the potential limitations imposed by the presence of a teacher model. Additionally, we improve the perceptual loss with edge-aware DISTS (EA-DISTS) to enhance the model’s ability to generate fine details. Our experiments demonstrate that, compared with previous diffusion-based methods requiring dozens or even hundreds of steps, our D3SR attains comparable or even superior results in both quantitative metrics and qualitative evaluations. Moreover, compared with other methods, D3SR achieves at least 3×3$\times$3 × faster inference speed and reduces parameters by at least 30%. We will release code and models at https://github.com/JianzeLi-114/D3SR.
@@ 1 Introduction
Real-world image super-resolution (Real-ISR) is a challenging task that aims to reconstruct high-resolution (HR) images from their low-resolution (LR) counterparts in real-world settings [46]. Most image super-resolution (SR) methods [20, 16, 21, 5, 6] use Bicubic downsampling of HR images to generate LR samples. These methods achieve good results in reconstructing simple degraded images. However, they struggle with the complex and unknown degradations widely existing in real-world scenarios. Previous research has predominantly employed generative adversarial networks (GANs) [10, 43, 62, 23] for Real-ISR task. In recent developments, diffusion models (DMs) [13, 36, 40] have emerged as a promising alternative. Recent Real-ISR methods have achieved outstanding performance by using diffusion models. In particular, some methods [52, 56, 25, 11, 60] leverage powerful pre-trained diffusion models, such as large-scale text-to-image (T2I) diffusion models like Stable Diffusion [36, 35]. These pre-trained T2I models provide extensive priors and powerful generative abilities. Most DM-based methods generate HR images by employing ControlNet models [64], conditioning on the LR inputs. However, these methods typically require tens to hundreds of diffusion steps to produce high-quality HR images. The introduction of ControlNet not only increases the number of model parameters but also further exacerbates inference latency. Consequently, DM-based multi-step diffusion methods often incur delays of tens of seconds when processing a single image, which significantly limits their practical application in real-world scenarios for low-level image reconstruction tasks.
@@
To accelerate the generation process of diffusion models, recent research has introduced numerous one-step diffusion methods [2, 39, 32, 37, 41, 58, 29]. These methods are known as diffusion distillation, which distill multi-step pre-trained diffusion models into one-step counterparts. Most of these approaches employ a knowledge distillation strategy, using the multi-step diffusion model as a teacher to train a one-step diffusion student model. These methods significantly reduce inference latency, and the quality of the generated images can be comparable to that of multi-step diffusion models. Real-ISR methods based on one-step diffusion models have become an increasingly popular research direction [34, 45, 54, 51]. These methods employ pre-trained multi-step diffusion models as teachers to guide the training process. They achieve promising results, with performance comparable to that of multi-step models. However, further performance improvement remains a challenging task. Existing methods rely on a multi-step teacher to conduct distillation. Nevertheless, this paradigm would inevitably come with several limitations. First, the multi-step teacher may hamper the effectiveness of distillation if the teacher itself is not strong enough. For example, variational score distillation (VSD) [47, 33, 51] is a one-step diffusion distillation method. It enhances the realism of generated images by optimizing the KL divergence between the scores of the teacher and student models. This approach heavily relies on the prior knowledge of the teacher diffusion model. The limitations of the multi-step diffusion model impose an upper bound on the performance of the aforementioned methods. Second, the widely used VSD approach does not exploit any HR data for training but merely depend on the prior knowledge embedded in the pre-trained parameters. If the distribution fitted by the multi-step diffusion model (teacher in VSD) deviates from the high-quality image distribution, it may lead to a loss of realism or the generation of fake textures in the student model’s images.
To overcome the aforementioned challenges, we propose a novel one-step diffusion model with a large-scale diffusion discriminator. Different from existing one-step diffusion models using distillation techniques with teacher models, we use a larger-scale diffusion model, SDXL [35], as a discriminator to leverage the powerful priors. Specifically, our diffusion discriminator aims to distill latent features with different noises from true data. At the same time, it enables us to perceive the true data distribution of high-quality super-resolution datasets. As a result, we overcome the performance limitations imposed by the teacher model’s upper bound. Additionally, we propose a simple and effective improvement to the perceptual loss, edge-aware DISTS (EA-DISTS), by capturing high-frequency details from extracted edges. Our comprehensive experiments indicate that D3SR achieves superior performance and less inference time among one-step DM-based Real-ISR models.
@@
  • We propose a simple but effective one-step diffusion distillation method, D3SR, which uses a large-scale diffusion model as a discriminator for adversarial training. Unlike previous one-step diffusion ISR methods, we do not use a pre-trained multi-step diffusion model as a teacher to guide the training. Instead, we employ a larger-scale diffusion model to guide the training process. This breaks through the performance limitations of teacher models in previous distillation methods.
  • We improve the perceptual loss by proposing the edge-aware DISTS (EA-DISTS) loss. Our EA-DISTS leverages image edges to enhance the model’s ability and improve the authenticity of reconstructed details.
[…]
2 Related Work
@@ 2.1 Real-World Image Super-Resolution
Real-world image super-resolution (Real-ISR) aims to recover high-resolution (HR) images from low-resolution (LR) observations in real-world scenarios. The complex and unknown degradation patterns in such scenarios make Real-ISR a challenging problem [14, 26, 15, 50]. To address this problem, models continuously evolve. Early image super-resolution models [20, 67, 66, 5, 6] typically rely on simple synthetic degradations like Bicubic downsampling for generating LR-HR pairs, resulting in subpar performance on real-world datasets. Later, GAN-based methods such as BSRGAN [62], Real-ESRGAN [43], and SwinIR-GAN [23] introduce more complex degradation processes. These methods achieve promising perceptual quality but encounter issues such as training instability. Additionally, they have limitations in preserving fine natural details. Recently, Stable Diffusion (SD) [36] is considered for addressing Real-ISR tasks due to its strong ability to capture complex data distributions and provide robust generative priors. Approaches such as StableSR [42], DiffBIR [25], and SeeSR [52] leverage pre-trained diffusion priors and ControlNet models [64] to enhance HR image generation. One-step diffusion models have gained widespread attention from researchers. Methods such as YONOS-SR [34], SinSR [45], OESEDiff [51], AddSR [54], TAD-SR [12], and TSD-SR [8] have achieved Real-ISR with diffusion models in a single sampling step.
@@ 2.2 Acceleration of Diffusion Models
Acceleration of diffusion models can reduce computational costs and inference time. Therefore, various strategies have been developed to enhance the efficiency of diffusion models in image generation tasks. Fast diffusion samplers [40, 17, 27, 30, 31, 68] have significantly reduced the number of sampling steps from 1,000 to 15$\sim$100 without requiring model retraining. However, further reducing the steps below 10 often leads to a performance drop. Under these circumstances, distillation techniques have made considerable progress in speeding up inference [2, 28, 32, 37, 41, 69, 58, 29, 9, 33]. For instance, Progressive Distillation (PD) methods [32, 37] have distilled pre-trained diffusion models to under 10 steps. Consistency models [41] have further reduced the steps to 2$\sim$4 with promising results. Instaflow [29] further achieves one-step generation through reflow [28] and distillation. Recent score distillation-based methods, such as Distribution Matching Distillation (DMD) [59, 57] and Variational Score Distillation (VSD) [49, 33], aim to achieve one-step text-to-image generation. They minimize the Kullback–Leibler (KL) divergence between the generated data distribution and the real data distribution. Although these approaches have made notable progress, they still face challenges, like high training costs and dependence on teacher models.
Figure 1: Training framework of D3SR. The left side represents the generator $\mathcal{G}_{\theta}$, which includes the pre-trained VAE and UNet from Stable Diffusion. Only the UNet is fine-tuned using LoRA, while other parameters remain frozen. The right side depicts the diffusion discriminator , which guides the training process without participating in inference. The discriminator extracts the UNet Mid-block outputs and processes them through an MLP to generate realism scores for different image regions. Both the downsample and middle blocks of the UNet in the discriminator are fine-tuned with LoRA, whereas the MLP is randomly initialized.
3 Method
In this section, we present our real-world image super-resolution (Real-ISR) model D3SR. First, in Sec. 3.1, we review the basics of diffusion models and introduce the D3SR generator. In Sec. 3.2, we introduce the diffusion distillation method using a large-scale diffusion discriminator. In Sec. 3.3, we introduce the edge-aware DISTS (EA-DISTS) perceptual loss. This loss improves texture details and enhances visual quality. Finally, in Sec. 3.4, we describe the training process for D3SR.
3.1 Preliminaries: Diffusion Models
@@
As illustrated in Fig. 1, we first employ the encoder $E_{\theta}$ to map the low-resolution (LR) image $x_{L}$ into the latent space, yielding $z_{L}$: $z_{L}=E_{\theta}(x_{L})$. Next, we perform a one denoising step to obtain the predicted noise $\hat{\epsilon}$ and compute the high-resolution (HR) latent representation $\hat{z}_{H}$:
@@
where $\epsilon_{\theta}$ denotes the denoising network parameterized by $\theta$, and $T_{L}$ is the diffusion time step. Unlike one-step text-to-image (T2I) diffusion models [41, 59], the input to the UNet of the Real-ISR diffusion models is not pure Gaussian noise. We set $T_{L}$ to an intermediate time step within the range [0, $T$], where $T$ is the total number of diffusion time steps. In Stable Diffusion (SD), $T=1,000$. Finally, we decode $\hat{z}_{H}$ using the decoder $D_{\theta}$ to reconstruct the HR image $\hat{x}_{H}$: $\hat{x}_{H}=D_{\theta}(\hat{z}_{H})$. The entire computation process of the generator can be expressed as $\hat{x}_{H}=\mathcal{G}_{\theta}(x_{L})$.
Figure 2: Visualization of features dimensionality reduction for the first 100 channels from the middle block outputs of the Stable Diffusion (SD) UNet. The distributions of the two types of image features are distinctly different.
Figure 3: Comparison of the performance of SD models with different scales as discriminators. As the model size increases, the performance of the generator improves accordingly.
3.2 Distillation with Large Diffusion Discriminator
Currently successful one-step diffusion distillation methods applied in image super-resolution include knowledge distillation (KD), variational score distillation (VSD), and others. Among them, VSD stands out due to its interesting principles and excellent performance. VSD works by training an additional diffusion network to fit the fake score $s_{\text{fake}}$ of the generated image, while using a frozen pre-trained diffusion model to obtain the real score $s_{\text{real}}$ based on its prior. The method then aligns the two scores’ differences using the Kullback-Leibler (KL) divergence, thus enhancing the realism of the generated image. However, in VSD, the absence of real image datasets implies that the upper bound of VSD is limited by the prior of the pre-trained diffusion model. The teacher model in VSD restricts the generative capacity of the student model, making further performance improvements challenging.
To address the issue incurred by a weak multi-step teacher, we propose to use a large-scale diffusion model to provide stronger guidance for one-step distillation. Figure 2 shows the distributions of real and generated images’ latent code at the middle block output of the UNet in the Stable Diffusion (SD) model. There is a clear difference in their distribution patterns. This suggests that using a pre-trained diffusion model as a discriminator has great potential. It can also avoid instability issues during the early stages of training. Figure 1 illustrates our training framework. We append an MLP block after the SD UNet middle block as a classifier to output the authenticity score for each patch.
During training, we input the forward diffusion results of both the generator’s predicted latent code $\hat{z}_{H}$, and the ground truth latent code $z_{H}=E_{\theta}(x_{H})$. The adversarial losses for updating the generator and discriminator are defined as: